# AI Agents vs. Agentic AI: A Conceptual Taxonomy, Applications and Challenges

### v3

@@ III-A Architectural Evolution: From AI Agents to Agentic AI Systems
Figure 8: Illustrating architectural evolution from traditional AI Agents to modern Agentic AI systems. It begins with core modules Perception, Reasoning, and Action and expands into advanced components including Specialized Agents, Advanced Reasoning & Planning, Persistent Memory, and Orchestration. The diagram further captures emergent properties such as Multi-Agent Collaboration, System Coordination, Shared Context, and Task Decomposition, all enclosed within a dotted boundary signifying layered modularity and the transition to distributed, adaptive agentic AI intelligence.
While both AI Agents and Agentic AI systems are grounded in modular design principles, Agentic AI significantly extends the foundational architecture to support more complex, distributed, and adaptive behaviors. As illustrated in Figure 8, the transition begins with core subsystems Perception, Reasoning, and Action, that define traditional AI Agents. Agentic AI enhances this base by integrating advanced components such as Specialized Agents, Advanced Reasoning & Planning, Persistent Memory, and Orchestration. The figure further emphasizes emergent capabilities including Multi-Agent Collaboration, System Coordination, Shared Context, and Task Decomposition, all encapsulated within a dotted boundary that signifies the shift toward reflective, decentralized, and goal-driven system architectures. This progression marks a fundamental inflection point in intelligent agent design. This section synthesizes findings from empirical frameworks such as LangChain [93], AutoGPT [94], and TaskMatrix [152], highlighting this progression in architectural sophistication.
@@ III-A1 Core Architectural Components of AI Agents
Foundational AI Agents are typically composed of four primary subsystems: perception, reasoning, action, and learning. These subsystems form a closed-loop operational cycle, commonly referred to as “Understand, Think, Act” from a user interface perspective, or “Input, Processing, Action, Learning” in systems design literature [153, 14].
  • Perception Module: This subsystem ingests input signals from users (e.g., natural language prompts) or external systems (e.g., APIs, file uploads, sensor streams). It is responsible for pre-processing data into a format interpretable by the agent’s reasoning module. For example, in LangChain-based agents [154, 93], the perception layer handles prompt templating, contextual wrapping, and retrieval augmentation via document chunking and embedding search.
  • Knowledge Representation and Reasoning (KRR) Module: At the core of the agent’s intelligence lies the KRR module, which applies symbolic, statistical, or hybrid logic to input data. Techniques include rule-based logic (e.g., if-then decision trees), deterministic workflow engines, and simple planning graphs. Reasoning in agents like AutoGPT [30] is enhanced with function-calling and prompt chaining to simulate thought processes (e.g., “step-by-step” prompts or intermediate tool invocations).
  • Action Selection and Execution Module: This module translates inferred decisions into external actions using an action library. These actions may include sending messages, updating databases, querying APIs, or producing structured outputs. Execution is often managed by middleware like LangChain’s “agent executor,” which links LLM outputs to tool calls and observes responses for subsequent steps [93].
@@ III-A2 Architectural Enhancements in Agentic AI
Agentic AI systems inherit the modularity of AI Agents but extend their architecture to support distributed intelligence, inter-agent communication, and recursive planning. The literature documents a number of critical architectural enhancements that differentiate Agentic AI from its predecessors [157, 158].
  • Ensemble of Specialized Agents: Rather than operating as a monolithic unit, Agentic AI systems consist of multiple agents, each assigned a specialized function e.g., a summarizer, a retriever, a planner. These agents interact via communication channels (e.g., message queues, blackboards, or shared memory). For instance, MetaGPT [151] exemplify this approach by modeling agents after corporate departments (e.g., CEO, CTO, engineer), where roles are modular, reusable, and role-bound.
  • Advanced Reasoning and Planning: Agentic systems embed recursive reasoning capabilities using frameworks such as ReAct [132], Chain-of-Thought (CoT) prompting [159], and Tree of Thoughts [160]. These mechanisms allow agents to break down a complex task into multiple reasoning stages, evaluate intermediate results, and re-plan actions dynamically. This enables the system to respond adaptively to uncertainty or partial failure.
  • Persistent Memory Architectures: Unlike traditional agents, Agentic AI incorporates memory subsystems to persist knowledge across task cycles or agent sessions [161, 162]. Memory types include episodic memory (task-specific history) [163, 164], semantic memory (long-term facts or structured data) [165, 166], and vector-based memory for retrieval-augmented generation (RAG) [167, 168]. For example, AutoGen [94] agents maintain scratchpads for intermediate computations, enabling stepwise task progression.
@@
The shift from isolated perception–reasoning–action loops to collaborative and reflective multi-agent workflows marks a key inflection point in the architectural design of intelligent systems. This progression positions Agentic AI as the next stage of AI infrastructure capable not only of executing predefined workflows but also of constructing, revising, and managing complex objectives across agents with minimal human supervision.
@@ IV Application of AI Agents and Agentic AI
To illustrate the real-world utility and operational divergence between AI Agents and Agentic AI systems, this study synthesizes a range of applications drawn from recent literature, as visualized in Figure 9. We systematically categorize and analyze application domains across two parallel tracks: conventional AI Agent systems and their more advanced Agentic AI counterparts. For AI Agents, four primary use cases are reviewed: (1) Customer Support Automation and Internal Enterprise Search, where single-agent models handle structured queries and response generation; (2) Email Filtering and Prioritization, where agents assist users in managing high-volume communication through classification heuristics; (3) Personalized Content Recommendation and Basic Data Reporting, where user behavior is analyzed for automated insights; and (4) Autonomous Scheduling Assistants, which interpret calendars and book tasks with minimal user input. In contrast, Agentic AI applications encompass broader and more dynamic capabilities, reviewed through four additional categories: (1) Multi-Agent Research Assistants that retrieve, synthesize, and draft scientific content collaboratively; (2) Intelligent Robotics Coordination, including drone and multi-robot systems in fields like agriculture and logistics; (3) Collaborative Medical Decision Support, involving diagnostic, treatment, and monitoring subsystems; and (4) Multi-Agent Game AI and Adaptive Workflow Automation, where decentralized agents interact strategically or handle complex task pipelines.
@@ IV-1 Application of AI Agents
    A practical example (Figure 10a) of this dual functionality can be seen in a multinational e-commerce company deploying an AI Agent-based customer support and internal search assistant. For customer support, the AI Agent integrates with the company’s CRM (e.g., Salesforce) and fulfillment APIs to resolve queries such as “Where is my order?” or “How can I return this item?”. Within milliseconds, the agent retrieves contextual data from shipping databases and policy repositories, then generates a personalized response using retrieval-augmented generation. For internal enterprise search, employees use the same system to query past meeting notes, sales presentations, or legal documents. When an HR manager types “summarize key benefits policy changes from last year,” the agent queries a Pinecone vector store embedded with enterprise documentation, ranks results by semantic similarity, and returns a concise summary along with source links. These capabilities not only reduce ticket volume and support overhead but also minimize time spent searching for institutional knowledge (like policies, procedures, or manuals). The result is a unified, responsive system that enhances both external service delivery and internal operational efficiency using modular AI Agent architectures.
@@
    Email Filtering and Prioritization: Within productivity tools, AI Agents automate email triage through content classification and prioritization. Integrated with systems like Microsoft Outlook and Superhuman, these agents analyze metadata and message semantics to detect urgency, extract tasks, and recommend replies. They apply user-tuned filtering rules, behavioral signals, and intent classification to reduce cognitive overload. Autonomous actions, such as auto-tagging or summarizing threads, enhance efficiency, while embedded feedback loops enable personalization through incremental learning [63].
@@
    A practical application of autonomous scheduling agents can be seen in corporate settings as depicted in Figure 10d where employees manage multiple overlapping responsibilities across global time zones. Consider an executive assistant AI agent integrated with Google Calendar and Slack that interprets a command like “Find a 45-minute window for a follow-up with the product team next week.” The agent parses the request, checks availability for all participants, accounts for time zone differences, and avoids meeting conflicts or working-hour violations. If it identifies a conflict with a previously scheduled task, it may autonomously propose alternative windows and notify affected attendees via Slack integration. Additionally, the agent learns from historical user preferences such as avoiding early Friday meetings and refines its suggestions over time. Tools like Reclaim AI and Clockwise exemplify this capability, offering calendar-aware automation that adapts to evolving workloads. Such assistants reduce coordination overhead, increase scheduling efficiency, and enable smoother team workflows by proactively resolving ambiguity and optimizing calendar utilization.
@@
OpenAI (2025) Deep Research OpenAI
@@
OpenAI (2025) Operator OpenAI
@@
Google (2025) Google Agentspace
@@
Amazon (2025) Amazon Nova
@@
Monica (2025) Manus Agenthttps://manus.im/
@@
Otter.ai (2025) Otter sales agent
@@
ClickUp (2025) ClickUp Brain
@@
Microsoft (2024) Microsoft Copilot
@@
Google DeepMind (2025) Project Astra
@@
Anthropic (2025) Claude 3.5 Sonnet
@@ IV-2 Appications of Agentic AI
    Multi-Agent Research Assistants: Agentic AI systems are increasingly deployed in academic and industrial research pipelines to automate multi-stage knowledge work. Platforms like AutoGen and CrewAI assign specialized roles to multiple agents retrievers, summarizers, synthesizers, and citation formatters under a central orchestrator. The orchestrator distributes tasks, manages role dependencies, and integrates outputs into coherent drafts or review summaries. Persistent memory allows for cross-agent context sharing and refinement over time. These systems are being used for literature reviews, grant preparation, and patent search pipelines, outperforming single-agent systems such as ChatGPT by enabling concurrent sub-task execution and long-context management [94].
@@
    Intelligent Robotics Coordination: In robotics and automation, Agentic AI underpins collaborative behavior in multi-robot systems. Each robot operates as a task specialized agent such as pickers, transporters, or mappers while an orchestrator supervises and adapts workflows. These architectures rely on shared spatial memory, real-time sensor fusion, and inter-agent synchronization for coordinated physical actions. Use cases include warehouse automation, drone-based orchard inspection, and robotic harvesting [151]. For instance, agricultural drone swarms may collectively map tree rows, identify diseased fruits, and initiate mechanical interventions. This dynamic allocation enables real-time reconfiguration and autonomy across agents facing uncertain or evolving environments.
    For example, in commercial apple orchards (Figure 11b), Agentic AI enables a coordinated multi-robot system to optimize the harvest season. Here, task-specialized robots such as autonomous pickers, fruit classifiers, transport bots, and drone mappers operate as agentic units under a central orchestrator. The mapping drones first survey the orchard and use vision-language models (VLMs) to generate high-resolution yield maps and identify ripe clusters. This spatial data is shared via a centralized memory layer accessible by all agents. Picker robots are assigned to high-density zones, guided by path-planning agents that optimize routes around obstacles and labor zones. Simultaneously, transport agents dynamically shuttle crates between pickers and storage, adjusting tasks in response to picker load levels and terrain changes. All agents communicate asynchronously through a shared protocol, and the orchestrator continuously adjusts task priorities based on weather forecasts or mechanical faults. If one picker fails, nearby units autonomously reallocate workload. This adaptive, memory-driven coordination exemplifies Agentic AI’s potential to reduce labor costs, increase harvest efficiency, and respond to uncertainties in complex agricultural environments far surpassing the rigid programming of legacy agricultural robots [151, 94].
@@
    Collaborative Medical Decision Support: In high-stakes clinical environments, Agentic AI enables distributed medical reasoning by assigning tasks such as diagnostics, vital monitoring, and treatment planning to specialized agents. For example, one agent may retrieve patient history, another validates findings against diagnostic guidelines, and a third proposes treatment options. These agents synchronize through shared memory and reasoning chains, ensuring coherent and safe recommendations. Applications include ICU management, radiology triage, and pandemic response. Real-world pilots show improved efficiency and decision accuracy compared to isolated expert systems [92].
@@
    Multi-Agent Game AI and Adaptive Workflow Automation: In simulation environments and enterprise systems, Agentic AI facilitates decentralized task execution and emergent coordination. Game platforms like AI Dungeon deploy independent NPC agents with goals, memory, and dynamic interactivity to create emergent narratives and social behavior. In enterprise workflows, systems such as MultiOn and Cognosys use agents to manage processes like legal review or incident escalation, where each step is governed by a specialized module. These architectures exhibit resilience, exception handling, and feedback-driven adaptability far beyond rule-based pipelines.
@@
Open Source (2023) GPT Engineer
@@
Salesforce (2024) Einstein Copilot
@@
Microsoft (2025) Github Agentic Copilot
@@
Atera (2025) Atera Agentic AI

### v4

@@ III-A Architectural Evolution: From AI Agents to Agentic AI Systems
Figure 8: Illustrating architectural evolution from traditional AI Agents to modern Agentic AI systems. It begins with core modules Perception, Reasoning and Action, and expands into advanced components including Specialized Agents, Advanced Reasoning & Planning, Persistent Memory, and Orchestration. The diagram further captures emergent properties such as Multi-Agent Collaboration, System Coordination, Shared Context, and Task Decomposition, all enclosed within a dotted boundary signifying layered modularity and the transition to distributed, adaptive agentic AI intelligence.
While both AI Agents and Agentic AI systems utilize modular design principles, Agentic AI significantly extends the foundational architecture to support more complex, distributed, and adaptive behaviors. As illustrated in Figure 8, the transition begins with core subsystems’ Perception, Reasoning, and Action, that define traditional AI Agents. Agentic AI enhances this foundation by integrating advanced components such as Specialized Agents, Advanced Reasoning & Planning, Persistent Memory, and Orchestration. The figure further emphasizes emergent capabilities including Multi-Agent Collaboration, System Coordination, Shared Context, and Task Decomposition, all encapsulated within a dotted boundary that signifies the shift toward proactive, decentralized, and goal-driven system architectures. As mentioned before, this progression marks a fundamental inflection point in intelligent agent design. This section synthesizes findings from empirical frameworks such as LangChain [99], AutoGPT [100], and TaskMatrix [158], highlighting this progression in architectural sophistication.
@@ III-A1 Core Architectural Components of AI Agents
Foundational AI Agents are typically composed of four primary subsystems: perception, reasoning, action, and learning. These subsystems form a closed-loop operational cycle, commonly referred to as “Understand, Think, Act, Learn” from a user interface perspective, or “Input, Processing, Action, Learning” in systems design literature [159, 21].
  • Perception Module: This subsystem intakes input signals from users (e.g., natural language prompts) or external systems (e.g., APIs, file uploads, sensor streams), and performs data pre-processing to create datasets in formats interpretable by the agent’s reasoning module. For example, in LangChain-based agents [160, 99], the perception layer handles prompt templating, contextual wrapping, and retrieval augmentation via document chunking and embedding search.
  • Knowledge Representation and Reasoning (KRR) Module: At the core of the agent’s intelligence lies the KRR module, which applies symbolic, statistical, or hybrid logic to input data. Techniques include rule-based logic (e.g., if-then decision trees), deterministic workflow engines, or simple planning graphs. Reasoning in agents like AutoGPT [37] is enhanced with function-calling and prompt chaining to simulate thought processes (e.g., “step-by-step” prompts or intermediate tool invocations).
  • Action Selection and Execution Module: This module translates inferred knowledge and decisions into external actions using an action library. These actions may include sending messages, updating databases, querying APIs, or producing structured outputs. Execution is often managed by middleware like LangChain’s “agent executor,” which links LLM outputs to tool calls and observes responses for subsequent steps [99].
@@ III-A2 Architectural Enhancements in Agentic AI
As discussed before, Agentic AI systems inherit the modularity of AI Agents but extend their architecture to support distributed intelligence, inter-agent communication, and iterative planning. The literature documents a number of critical architectural enhancements that differentiate Agentic AI from its predecessors and enable them to be highly versatile and adaptive [163, 164].
  • Ensemble of Specialized Agents: Rather than operating as a monolithic unit, Agentic AI systems consist of multiple agents, each assigned a specialized function or task (e.g., a summarizer, a retriever, or a planner). These agents interact via communication channels (e.g., message queues, blackboards, or shared memory). For instance, MetaGPT [157] highlights this approach by modeling agents after corporate departments (e.g., CEO, CTO, engineer), where roles are modular, reusable, and role-bound.In this context, ”role-bound” means that each agent’s behavior and responsibilities are strictly defined by its assigned role, limiting its scope of action to that specific functional domain.
  • Advanced Reasoning and Planning: Agentic systems embed iterative reasoning capabilities using frameworks such as ReAct [138], Chain-of-Thought (CoT) prompting [165], and Tree of Thoughts [166]. These mechanisms allow agents to break down a complex task into multiple reasoning stages, evaluate intermediate results, and re-plan actions dynamically. This enables the system to respond adaptively to uncertainty or partial failure.
  • Persistent Memory Architectures: Unlike traditional agents, Agentic AI incorporates memory subsystems to preserve and persist knowledge across task cycles or agent sessions [167, 168]. Memory types include episodic memory (task-specific history) [169, 170], semantic memory (long-term facts or structured data) [171, 172], and vector-based memory for RAG [173, 174]. For example, AutoGen [100] agents maintain scratchpads for intermediate computations, enabling stepwise task progression.
@@
The shift from isolated perception–reasoning–action loops to collaborative and self-evaluative multi-agent workflows marks a key turning point in the architectural design of intelligent systems, enabling agents not only to act but also to reflect, learn, and improve over time [175]. This progression positions Agentic AI as the next stage of AI infrastructure capable not only of executing predefined workflows but also of constructing, revising, and managing complex objectives across agents with minimal human supervision.
@@ IV Application of AI Agents and Agentic AI
To illustrate the real-world utility and operational divergence between AI Agents and Agentic AI systems, this study synthesizes a range of applications drawn from recent literature, as visualized in Figure 9. We systematically categorize and analyze application domains across two parallel tracks: conventional AI Agent systems and their more advanced Agentic AI counterparts. For AI Agents, four primary use cases are reviewed: (1) Customer Support Automation and Internal Enterprise Search, where single-agent models handle structured queries and response generation; (2) Email Filtering and Prioritization, where agents assist users in managing high-volume communication through classification heuristics; (3) Personalized Content Recommendation and Basic Data Reporting, where user behavior is analyzed for automated insights; and (4) Autonomous Scheduling Assistants, which interpret calendars and book tasks with minimal user input. In contrast, Agentic AI applications encompass broader and more dynamic capabilities, reviewed and discussed in four categories as well: (1) Multi-Agent Research Assistants that retrieve, synthesize, and draft scientific content collaboratively; (2) Intelligent Robotics Coordination, including drone and multi-robot systems in fields like agriculture and logistics; (3) Collaborative Medical Decision Support, involving diagnostic, treatment, and monitoring subsystems; and (4) Multi-Agent Game AI and Adaptive Workflow Automation, where decentralized agents interact strategically or handle complex task pipelines.
@@ IV-1 Application of AI Agents
    A practical example (Figure 10a) of this dual functionality can be seen in a multi-national e-commerce company deploying an AI Agent-based customer support and internal search assistant. For customer support, the AI Agent integrates with the company’s Customer Relationship Management (CRM) system (e.g., Salesforce) and fulfillment APIs to resolve queries such as “Where is my order?” or “How can I return this item?”. Within milliseconds, the agent retrieves contextual data from shipping databases and policy repositories, then generates a personalized response using retrieval-augmented generation. For internal enterprise search, employees use the same system to query past meeting notes, sales presentations, or legal documents. When an HR manager types “summarize key benefits of policy changes from last year,” the agent queries a Pinecone vector store embedded with enterprise documentation, ranks results by semantic similarity, and returns a concise summary along with source links. These capabilities not only reduce ticket volume and support overhead but also minimize time spent searching for institutional knowledge (like policies, procedures, or manuals). The result is a unified, responsive system that enhances both external service delivery and internal operational efficiency using modular AI Agent architectures.
@@
    Email Filtering and Prioritization: As one of the important productivity tools, AI Agents automate email triage through content classification and prioritization. Integrated with systems like Microsoft Outlook and Superhuman, these agents analyze metadata and message semantics to detect urgency, extract tasks, and recommend replies. They apply user-tuned filtering rules, behavioral signals, and intent classification to reduce cognitive overload. Autonomous actions, such as auto-tagging or summarizing threads, enhance efficiency, while embedded feedback loops enable personalization through incremental learning [66].
@@
    A practical application of autonomous scheduling agents can be seen in corporate settings as depicted in Figure 10d where employees manage multiple overlapping responsibilities across global time zones. Consider an executive assistant AI agent integrated with Google Calendar and Slack that interprets a command like “Find a 45-minute window for a follow-up with the product team next week.” The agent parses the request, checks availability for all participants, accounts for time zone differences, and avoids meeting conflicts or working-hour violations. If it identifies a conflict with a previously scheduled task, it may autonomously propose alternative windows and notify affected attendees via Slack integration. Additionally, the agent learns from historical user preferences such as avoiding early Friday meetings and refines its suggestions over time. Tools like Reclaim AI and Clockwise further illustrate this capability, offering calendar-aware automation that adapts to evolving workloads. Such assistants reduce coordination overhead, increase scheduling efficiency, and enable smoother team workflows by proactively resolving ambiguity and optimizing calendar utilization.
@@
OpenAI (2025) Source Link
@@
OpenAI (2025) Source Link
@@
Google (2025) Source Link
@@
Amazon (2025) Source Link
@@
Monica (2025) Source Linkhttps://manus.im/
@@
Otter.ai (2025) Source Link
@@
ClickUp (2025) Source Link
@@
Microsoft (2024) Source Link
@@
Google DeepMind (2025) Source Link
@@
Anthropic (2025) Source Link
@@ IV-2 Appications of Agentic AI
    Multi-Agent Research Assistants: Agentic AI systems are increasingly deployed in academic and industrial research pipelines to automate multi-stage knowledge compilation. Platforms like AutoGen and CrewAI assign specialized roles to multiple agent retrievers, summarizers, synthesizers, and citation formatters under a central orchestrator. The orchestrator distributes tasks, manages role dependencies, and integrates outputs into coherent drafts or review summaries. Persistent memory allows for cross-agent context sharing and refinement over time. These systems are being used for literature reviews, grant preparation, and patent search pipelines, outperforming single-agent systems such as ChatGPT by enabling concurrent sub-task execution and long-context management [100].
@@
    Intelligent Robotics Coordination: In robotics and automation, Agentic AI enable collaborative behavior in multi-robot systems. Each robot operates as a task specialized agent such as pickers, transporters, or mappers while an orchestrator supervises and adapts workflows. These architectures rely on shared spatial memory, real-time sensor fusion, and inter-agent synchronization for coordinated physical actions. Use cases include warehouse automation, drone-based orchard inspection, and robotic harvesting [157]. For instance, a swarm of agricultural drones may collectively map tree rows, identify diseased fruits, and initiate mechanical interventions. This dynamic allocation enables real-time reconfiguration and autonomy across agents facing uncertain or evolving environments.
    For example, in commercial apple orchards (Figure 11b), Agentic AI enables a coordinated multi-robot system to optimize fruit harvesting. Here, task-specialized robots such as autonomous pickers, fruit classifiers, transport bots, and drone mappers operate as agentic units under a central orchestrator. The mapping drones first survey the orchard and use vision-language models (VLMs) to generate high-resolution yield maps and identify ripe fruit clusters. This spatial data is shared via a centralized memory layer accessible by all agents. Picker robots are then assigned to high-density zones, guided by path-planning agents that optimize routes around obstacles and labor zones. Simultaneously, transport agents dynamically haul fruit containers or bins between pickers and storage, adjusting tasks in response to picker load levels and terrain changes. All agents communicate asynchronously through a shared protocol, and the coordinator continuously adjusts task priorities based on weather forecasts or mechanical faults. If one picker fails, nearby units autonomously reallocate workload. This adaptive, memory-driven coordination exemplifies Agentic AI’s potential to reduce labor costs, increase harvest efficiency, and respond to uncertainties in complex agricultural environments surpassing the rigid programming of conventional agricultural robots [157, 100].
@@
    Collaborative Medical Decision Support: In high-stakes clinical environments, Agentic AI enables distributed medical reasoning by assigning tasks such as diagnostics, vital monitoring, and treatment planning to specialized agents. For example, one agent may retrieve patient history, another validates findings against diagnostic guidelines, and a third proposes treatment options (as seen in China’s world’s first Agentic AI hospital [176]). These agents synchronize through shared memory and reasoning chains, ensuring coherent and safe recommendations. Applications include ICU management, radiology triage, and pandemic response. Although real-world implementations are still lacking due to the nascent nature of the field, studies support the potential of Agentic AI to revolutionize the healthcare sector [177].
@@
    Multi-Agent Game AI and Adaptive Workflow Automation: In simulation environments and enterprise systems, Agentic AI systems facilitate decentralized task execution and effective coordination. Game platforms like AI Dungeon deploy independent NPC agents with goals, memory, and dynamic interactivity to create emergent narratives and social behavior. In enterprise workflows, systems such as MultiOn and Cognosys use agents to manage processes like legal review or incident escalation, where each step is governed by a specialized module. These architectures exhibit resilience, exception handling, and feedback-driven adaptability far beyond rule-based pipelines [178].
@@
Open Source (2023) Source Link
@@
Salesforce (2024) Source Link
@@
Microsoft (2025) Source Link
@@
Atera (2025) Source Link